# VesselSAM: Leveraging SAM for Aortic Vessel Segmentation with LoRA and Atrous Attention

### v1

@@ Abstract
Medical image segmentation is crucial for clinical diagnosis and treatment planning, particularly for complex anatomical structures like vessels. In this work, we propose VesselSAM, a modified version of the Segmentation Anything Model (SAM), specifically designed for aortic vessel segmentation. VesselSAM incorporates AtrousLoRA, a novel module that combines Atrous Attention with Low-Rank Adaptation (LoRA), to improve segmentation performance. Atrous Attention enables the model to capture multi-scale contextual information, preserving both fine local details and broader global context. At the same time, LoRA facilitates efficient fine-tuning of the frozen SAM image encoder, reducing the number of trainable parameters and ensuring computational efficiency. We evaluate VesselSAM on two challenging datasets: the Aortic Vessel Tree (AVT) dataset and the Type-B Aortic Dissection (TBAD) dataset. VesselSAM achieves state-of-the-art performance with DSC scores of 93.50%, 93.25%, 93.02%, and 93.26% across multiple medical centers. Our results demonstrate that VesselSAM delivers high segmentation accuracy while significantly reducing computational overhead compared to existing large-scale models. This development paves the way for enhanced AI-based aortic vessel segmentation in clinical environments. The code and models will be released at https://github.com/Adnan-CAS/AtrousLora.
@@ I Introduction
Medical imaging stands at the cutting edge of modern healthcare, serving a vital tool in diagnosing and treating various diseases. Within this domain, medical image segmentation is a critical component aiming to delineate structures such as organs, tumors, and vessels [1]. Aortic vessel segmentation is particularly important for diagnosing cardiovascular diseases, enabling precise assessments of vascular health and supporting interventions such as stent placement and aneurysm monitoring. It plays a crucial role in computer-aided diagnosis, treatment planning, and surgical interventions [2]. With the rapid advancement of computational resources and the growing availability of medical data, Vision Transformers (ViTs) have emerged as a transformative approach in medical image analysis [3]. Unlike traditional convolutional models, ViTs utilize self-attention mechanisms to capture long-range dependencies and global context [4], significantly improving the ability to model complex structures within medical images [5].
This paradigm shift has paved the way for more advanced segmentation techniques like the Segmentation Anything Model (SAM) [6], Swin-Unet [7], UNETR [8], SAMMedAI [9], and MedSAM [10] which leverage the power of ViT’s for accurate and efficient segmentation tasks. SAM enables users to generate segmentation masks using interactive prompts like clicks, bounding boxes and text. Its remarkable zero-shot and few-shot abilities have proven highly effective for natural images, gaining wide attention. However, while SAM excels in natural image segmentation, recent studies have highlighted its limitations in the medical domain[11][12].
@@
Medical images, often characterized by low contrast, ambiguous tissue boundaries, and small regions of interest, present unique challenges to SAM [13]. Several recent approaches [14, 15, 16] have sought to fine-tune SAM for medical image segmentation by incorporating domain-specific enhancements. However, fine-tuning these models requires significant computational resources, due to the extensive number of parameters in foundation models like SAM. Furthermore, training large models with limited task-specific data often results in overfitting and inferior performance. To overcome these challenges, parameter-efficient fine-tuning (PEFT) approaches such as Low-Rank Adaptation (LoRA) [17] have been developed as a viable solution. Various methods have incorporated LoRA with SAM to address the computational efficiency while maintaining performance gains, specifically for medical image segmentation[18][19].
@@
Despite these efforts, several fundamental intrinsic limitations of SAM persist. SAM’s image encoder, based on plain ViTs, inherently lacks crucial vision-specific inductive biases needed to capture local patterns and fine details essential for dense predictions in medical imaging [20]. Additionally, SAM’s ViT-based architecture relies on global attention without integrating regional attention or sparse attention mechanisms, which are vital for focusing on relevant regions and reducing computational overhead [21]. While regional attention, helps capture spatial hierarchies at different scales, SAM’s reliance on global attention limits its ability to focus on smaller, intricate regions in medical images. On the contrary, the absence of sparse attention prevents SAM from efficiently capturing global context without a significant computational cost. These limitations make SAM prone to errors, such as hallucinating small, disconnected components in its segmentation [4][9], especially when modeling structures like vessels, tumors, or lesions. To enhance the performance of plain ViTs in dense prediction tasks, recent research has combined Transformer and convolutional features [22] [23]. Work [24] integrates atrous attention with ViTs, enabling multi-scale feature extraction while preserving resolution. Atrous attention combines regional and sparse attention, allowing the model to focus on local details while still capturing the broader context.
@@
Inspired by the work[24], we propose VesselSAM, a model that integrates Atrous Attention with SAM, leveraging both global transformer attention and local convolutional inductive biases. VesselSAM incorporates several key innovations to enhance SAM’s capabilities. First, we integrate Atrous Spatial Pyramid Pooling (ASPP) to capture multi-scale contextual information, enabling the model to handle both small and large anatomical structures without sacrificing spatial resolution[22]. Additionally, Atrous Attention mechanisms are introduced, combining dilated windows at different scales to balance local feature extraction with global contextual understanding, allowing the model to focus on fine details while maintaining a comprehensive view of the entire image [23]. Furthermore, VesselSAM incorporates LoRA[18] layers to fine-tune the model efficiently, reducing the need for computationally expensive full retraining while ensuring high performance across various medical segmentation tasks.
@@
  • We propose VesselSAM, a novel model that integrates AtrousLoRA,a module that combines Atrous Attention within Low-Rank Adaptation (LoRA) to enhance the SAM architecture for vascular image segmentation, particularly aortic vessel segmentation. AtrousLoRA enables VesselSAM to capture both local and global features efficiently, improving segmentation accuracy while keeping the pre-trained image encoder frozen. The integration of Atrous Attention allows for multi-scale feature extraction through dilated convolutions, while LoRA reduces the number of trainable parameters without compromising model performance.
  • AtrousLoRA is integrated into VesselSAM and comprises two key modules: the Atrous Spatial Pyramid Pooling (ASPP) module and the Attention mechanism. The ASPP module uses dilated convolutions at different rates to capture multi-scale contextual information, enabling VesselSAM to focus on both fine details, such as small vessel boundaries, and broader anatomical structures without losing spatial resolution. Meanwhile, the Attention mechanism balances local feature extraction with global context, guiding VesselSAM to focus on relevant anatomical regions and improving segmentation performance.
  • AtrousLoRA leverages LoRA’s key concept of applying a low-rank constraint to the transformer features, enabling VesselSAM to fine-tune the model efficiently with only 7% of the trainable parameters. This reduces the computational cost significantly, making VesselSAM more suitable for tasks with limited data, as it eliminates the need for full retraining while maintaining high performance.
  • We evaluate VesselSAM on multiple challenging benchmark datasets, including the Aortic Vessel Tree (AVT) Segmentation dataset and the imageTBAD dataset. The results show that VesselSAM, with AtrousLoRA, outperforms baseline methods in segmentation accuracy, robustness, and computational efficiency, particularly for aortic vessel segmentation.
The rest of this article is organized as follows. Section II provides an overview of related work, focusing on medical foundation models, parameter-efficient fine-tuning strategies, and the application of atrous convolution in Vision Transformers (ViTs). Section III details the methodology, beginning with a preliminary description of the SAM architecture, followed by the introduction of VesselSAM, LoRA, and AtrousLoRA, as well as the integration of the Atrous Attention Module for capturing both local and global features. Section III further discusses the role of the Prompt Encoder and Mask Decoder in generating high-resolution segmentation masks. Section IV outlines the experimental setup, including dataset descriptions, evaluation metrics, and presents both quantitative and qualitative results. It also includes an ablation study, addressing the model’s limitations and suggesting potential future directions. Finally, Section V concludes the article.
Figure 1: The architecture of the proposed VesselSAM framework. The model combines frozen pre-trained transformer blocks with learnable LoRA-Atrous Attention layers, enabling enhanced multi-scale feature extraction through Atrous Spatial Pyramid Pooling (ASPP) and attention mechanisms. Image embeddings generated by the frozen image encoder are fused with a bounding box prompt from the frozen prompt encoder and processed by the learnable mask decoder to produce the final segmentation. Blue represents frozen components, while orange denotes learnable parameters.
@@ II-A ViT and SAM Based Medical Foundation Models
Vision Transformers (ViTs) based medical foundation models have significantly impacted medical image segmentation, with models like UNETR[8] leading the way. UNETR utilizes a ViT-based encoder to effectively capture global context while integrating it with a U-Net architecture for effective medical image segmentation. In contrast, SAM-based medical foundation models leveraging transformer architectures have demonstrated impressive performance across a broad range of natural image segmentation tasks. However, their application in the medical domain has faced limitations due to the unique challenges of medical images, such as low contrast and complex anatomical structures. Recognizing these limitations, MedSAM[10] sought to improve SAM’s performance in medical image segmentation by freezing the large pre-trained image encoder and prompt encoder, while fine-tuning only the lightweight mask decoder on domain-specific medical datasets. This approach leverages SAM’s powerful pre-trained architecture while adapting its mask prediction capabilities to the medical domain.
@@ II-B Parameter-Efficient Model Fine-Tuning
The concept of Parameter-Efficient Fine-Tuning (PEFT) has emerged as an effective way to adapt large foundational models like SAM to specific downstream tasks with minimal additional parameter costs. One prominent PEFT approach, LoRA (Low-Rank Adaptation), has been successfully incorporated into SAM-based models. For example, SAMed[11] applied LoRA to SAM’s frozen image encoder, fine-tuning the LoRA layers, the prompt encoder, and the mask decoder together on medical datasets like Synapse multiorgan, demonstrating significant performance improvements. Similarly, SAMAdp[19] introduces a lightweight adapter module to enhance SAM’s performance in challenging segmentation tasks. By integrating task-specific prompts and adapters, it improves accuracy while maintaining computational efficiency, demonstrating adaptability across various domains. Other works have taken different approaches to enhancing SAM for medical applications. SAMMed[9] evaluated SAM across 53 public medical imaging datasets, revealing that while SAM has strong zero-shot segmentation capabilities but it often underperforms without fine-tuning.
@@ II-C Atrous Convolution in ViTs
The use of Atrous Convolution (dilated convolution) in ViTs has gained attention as a powerful method to capture both local priors and global context, essential for segmentation tasks[21]. Atrous convolution increases the receptive field by ”skipping” pixels, allowing the model to capture information over larger regions without downsampling, thus preserving fine-grained details while enhancing the ability to model broader spatial relationships. This technique, initially popularized in DeepLab [6] for convolutional networks, has proven effective in extracting multi-scale features, which are crucial for complex segmentation tasks involving varying object sizes. In ViTs, where image features are typically processed as non-overlapping patches, integrating Atrous Convolutions enables the model to learn hierarchical spatial dependencies. By applying dilated convolutions at multiple rates, Atrous Spatial Pyramid Pooling (ASPP) modules allow the model to capture multi-scale contextual information, bridging the gap between local interactions and global dependencies. This approach is particularly beneficial in tasks requiring detailed segmentation, where capturing both local fine details and global context is necessary for accurate predictions. Recent advancements have shown that Atrous Convolutions are crucial for improving the performance of ViTs in segmentation tasks, particularly in domains such as medical imaging. In our model, we leverage the power of ASPP and Attention mechanisms to enhance the ViT encoder’s ability to capture both local priors and global context, effectively enabling the model to handle complex, high-resolution segmentation tasks with greater accuracy.
Figure 2: LoRA and AtrousLoRA: A comparative overview with detailed insights into the Atrous Attention Module. Both LoRA and Atrous-LoRA introduce a trainable encoder-decoder structure that operates in parallel with frozen pre-trained weights. (a) LoRA applies a low-rank constraint on the weight updates by factorizing them into smaller matrices (b) AtrousLoRA extends this approach by incorporating Atrous Attention Module into the bottleneck of LoRA, leveraging multi-scale dilated convolution operations for enhanced feature extraction. (c) The Atrous Attention Module features an Atrous Spatial Pyramid Pooling (ASPP) module with various dilation rates, global image-level pooling, and an attention mechanism that refines feature maps through element-wise multiplication with attention weights.
@@ III-A Overview
VesselSAM is a promptable segmentation model designed to enhance vascular structure segmentation in medical images. Building upon the standard SAM (Segment Anything Model) framework, VesselSAM integrates key modifications such as the Atrous Attention Module and LoRA (Low-Rank Adaptation) layers to improve performance for medical image segmentation. The image encoder and prompt encoder from the SAM model are frozen to preserve their pre-trained features, while the Atrous Attention module and LoRA layers enhance the model’s ability to capture multi-scale features and optimize training efficiency. The final segmentation is generated by a mask decoder that refines the fused embeddings using cross-attention mechanisms. The overall design ensures VesselSAM is a robust and efficient model for medical image segmentation tasks, particularly vascular segmentation in aortic imaging.

### v2

# xcolorIncompatible color definition

@@ Abstract
Medical image segmentation is crucial for clinical diagnosis and treatment planning, especially when dealing with complex anatomical structures such as vessels. However, accurately segmenting vessels remains challenging due to their small size, intricate edge structures, and susceptibility to artifacts and imaging noise. In this work, we propose VesselSAM, an enhanced version of the Segment Anything Model (SAM), specifically tailored for aortic vessel segmentation. VesselSAM incorporates AtrousLoRA, a novel module integrating Atrous Attention and Low-Rank Adaptation (LoRA), to enhance segmentation performance. Atrous Attention enables the model to capture multi-scale contextual information, preserving both fine-grained local details and broader global context. Additionally, LoRA facilitates efficient fine-tuning of the frozen SAM image encoder, reducing the number of trainable parameters and thereby enhancing computational efficiency. We evaluate VesselSAM using two challenging datasets: the Aortic Vessel Tree (AVT) dataset and the Type-B Aortic Dissection (TBAD) dataset. VesselSAM achieves state-of-the-art performance, attaining DSC scores of 93.50%, 93.25%, 93.02%, and 93.26% across multi-center datasets. Our results demonstrate that VesselSAM delivers high segmentation accuracy while significantly reducing computational overhead compared to existing large-scale models. This development paves the way for enhanced AI-based aortic vessel segmentation in clinical environments. The code and models will be released at https://github.com/Adnan-CAS/AtrousLora.
@@ I Introduction
Medical imaging stands at the cutting edge of modern healthcare, serving a vital tool in diagnosing and treating various diseases. Within this domain, medical image segmentation is a critical component aiming to delineate structures such as organs, tumors, and vessels [1]. Aortic vessel segmentation is crucial for diagnosing cardiovascular diseases, enabling precise vascular health assessments and facilitating interventions such as stent placement and aneurysm monitoring. It plays a crucial role in computer-aided diagnosis, treatment planning, and surgical interventions [2]. With the rapid advancements in computational resources and the increasing availability of medical data, Vision Transformers (ViTs) have emerged as a revolutionary approach in medical image analysis [3]. Unlike traditional convolutional models, ViTs employ self-attention mechanisms to capture long-range dependencies and global context [4], significantly enhancing their ability to model complex structures within medical images [5].
This paradigm shift has led to the development of advanced segmentation techniques, such as Segment Anything Model (SAM) [6], Swin-Unet [7], UNETR [8], SAMMed [9], and MedSAM [10], which leverage the power of ViT’s for highly accurate and computationally efficient segmentation tasks. The SAM enables users to generate segmentation masks through interactive prompts, such as clicks, bounding boxes, and text. Its exceptional zero-shot and few-shot capabilities have demonstrated strong effectiveness in natural image segmentation, garnering significant attention. However, despite the SAM’s success in natural image segmentation, recent studies have identified several limitations in its application to medical imaging [11][12]. These challenges stem from the inherent differences between natural and medical imaging data. Medical imaging datasets typically exhibit low contrast, ambiguous tissue boundaries, and small regions of interest. These limitations hinder SAM’s ability to generalize effectively without further fine-tuning [13].
@@
Recent studies [14, 15, 16] have sought to fine-tune SAM for medical image segmentation by incorporating domain-specific enhancements. However, fine-tuning these models demands substantial computational resources due to the large number of parameters in foundation models like SAM. Moreover, training large models on limited task-specific data frequently leads to overfitting and suboptimal performance. To address these challenges, parameter-efficient fine-tuning (PEFT) methods, such as Low-Rank Adaptation (LoRA) [17] have emerged as promising solutions. Several techniques have integrated LoRA into SAM to enhance computational efficiency while preserving performance, particularly in medical image segmentation [18][19].
@@
Despite these advancements, several fundamental intrinsic limitations of SAM persist. The SAM’s image encoder, based on plain ViTs, inherently lacks crucial vision-specific inductive biases needed to capture local patterns and fine details essential for dense predictions in medical imaging [20]. Additionally, SAM’s ViT-based architecture relies on global attention without integrating regional attention or sparse attention mechanisms, which are vital for focusing on relevant regions and reducing computational overhead [21]. Although regional attention captures spatial hierarchies at multiple scales, SAM’s reliance on global attention restricts its ability to focus on smaller, intricate regions in medical images. Moreover, the lack of sparse attention inhibits SAM from effectively capturing global context without incurring substantial computational costs. These limitations render SAM susceptible to errors, including the hallucination of small, disconnected components in segmentation [4][9], particularly when modeling structures such as vessels, tumors, or lesions. To enhance the performance of plain ViTs in dense prediction tasks, recent research has combined Transformer and convolutional features [22][23]. A study [24] integrates atrous attention with ViTs, enabling multi-scale feature extraction while preserving spatial resolution. Atrous attention unifies regional and sparse attention, enabling the model to focus on local details while simultaneously capturing the broader context.
@@
Inspired by the work[24], we propose VesselSAM, a model that integrates Atrous Attention with SAM, leveraging both global attention and local convolutional inductive biases. VesselSAM incorporates several key innovations to enhance SAM’s capabilities. First, we incorporate Atrous Spatial Pyramid Pooling (ASPP) to capture multi-scale contextual information, enabling the model to handle both small and large anatomical structures without sacrificing spatial resolution[22]. Additionally, Atrous Attention mechanisms are introduced, combining dilated windows at different scales to balance local feature extraction with global contextual understanding, allowing the model to focus on fine details while maintaining a comprehensive view of the entire image [23]. Furthermore, VesselSAM incorporates LoRA[18] layers to fine-tune the model efficiently, reducing the need for computationally expensive full retraining while ensuring high performance across diverse medical segmentation tasks.
@@
  • We propose VesselSAM, a novel segmentation model that integrates AtrousLoRA, a module designed to enhance the Segment Anything Model (SAM) for vascular image segmentation, particularly for aortic vessel segmentation. The AtrousLoRA enables efficient feature extraction by capturing both local and global information, improving segmentation accuracy while keeping the pre-trained image encoder frozen.
  • We introduce AtrousLoRA as an extension of LoRA, incorporating Atrous Attention Module to improve multi-scale feature extraction while reducing trainable parameters. AtrousLoRA integrates the Atrous Spatial Pyramid Pooling (ASPP) module and an Attention mechanism. The ASPP module utilizes dilated convolutions at different rates to capture multi-scale contextual information, allowing VesselSAM to process both fine details and broader anatomical structures without compromising spatial resolution. Meanwhile, the Attention mechanism balances local feature extraction with global context, enhancing the model’s ability to focus on clinically relevant anatomical regions.
  • We develop a parameter efficient fine-tuning (PEFT) strategy using AtrousLoRA, enabling the model to achieve high segmentation accuracy with only 7% of the trainable parameters. This approach drastically lowers computational costs and makes VesselSAM highly adaptable to medical segmentation tasks, even with limited available data.
  • We evaluate VesselSAM on multiple benchmark datasets, including the Aortic Vessel Tree (AVT) Segmentation dataset and the imageTBAD dataset. Experimental results demonstrate that VesselSAM consistently outperforms existing baseline methods in terms of segmentation accuracy, robustness, and computational efficiency, particularly for aortic vessel segmentation.
Figure 1: The architecture of the proposed VesselSAM framework. The model combines frozen pre-trained transformer blocks with learnable AtrousLoRA layers, enabling enhanced multi-scale feature extraction through Atrous Attention Module. Image embeddings generated by the frozen image encoder are fused with a bounding box prompt from the frozen prompt encoder and processed by the learnable mask decoder to produce the final segmentation. Blue represents frozen components, while orange denotes learnable parameters.
@@ II-A ViT and SAM Based Medical Foundation Models
Vision Transformers (ViTs) based medical foundation models have significantly impacted medical image segmentation, with models like UNETR[8] leading the way. UNETR employs a ViT-based encoder to effectively capture global context while integrating it with a U-Net architecture for precise medical image segmentation. In contrast, SAM-based medical foundation models, which leverage transformer architectures, have exhibited impressive performance across natural image segmentation tasks. However, their direct application to medical image segmentation remains challenging due to unique domain-specific constraints, such as low contrast, complex anatomical structures, and limited labeled data. Recognizing these limitations, MedSAM [10] sought to enhance SAM’s segmentation performance in the medical domain by freezing the pre-trained image encoder and prompt encoder, while fine-tuning only the lightweight mask decoder on domain-specific medical datasets. This approach effectively leverages SAM’s large-scale pre-trained features while adapting its mask prediction capabilities to medical imaging domain.
@@ II-B Parameter-Efficient Model Fine-Tuning
The concept of Parameter-Efficient Fine-Tuning (PEFT) has emerged as an effective strategy to adapt large foundational models like SAM to specific downstream tasks with minimal additional parameter costs. One prominent PEFT approach, LoRA (Low-Rank Adaptation), has been successfully incorporated into SAM-based models. For instance,, SAMed[11] applied LoRA to SAM’s frozen image encoder, fine-tuning the LoRA layers, the prompt encoder, and the mask decoder together on medical datasets like Synapse multiorgan, demonstrating significant performance improvements. Similarly, SAMAdp[19] introduced a lightweight adapter module to enhance SAM’s segmentation performance in challenging tasks. By integrating task-specific prompts and adapters, SAMAdp improves segmentation accuracy while maintaining computational efficiency, demonstrating broad adaptability across diverse domains. Other works have pursued different approaches to optimize SAM for medical imaging applications. SAMMed[9] systematically evaluated SAM across 53 public medical imaging datasets, revealing that while SAM demonstrates strong zero-shot segmentation capabilities, its performance often degrades without fine-tuning, reinforcing the need for domain-specific adaptation.
@@ II-C Atrous Convolution in ViTs
Atrous Convolution (dilated convolution) has emerged as a powerful technique in Vision Transformers (ViTs) to enhance both local feature extraction and global contextual modeling, which are critical for segmentation tasks[25][26][27]. Atrous convolution expands the receptive field by introducing pixel “skipping”, enabling the model to capture multi-scale spatial dependencies without downsampling. This preserves fine-grained details while improving the ability to model broader spatial relationships. Initially introduced in DeepLab [20] for convolutional networks, Atrous Convolution has proven highly effective in extracting multi-scale features, which is crucial for handling segmentation tasks involving objects of varying sizes. In ViTs, where image features are typically processed as non-overlapping patches, integrating Atrous Convolutions enhances the model’s ability to learn hierarchical spatial dependencies. Specifically, Atrous Spatial Pyramid Pooling (ASPP) modules apply dilated convolutions at multiple rates, allowing the model to capture multi-scale contextual information [28], bridging the gap between local interactions and global dependencies. This approach is particularly beneficial in tasks requiring detailed segmentation, where capturing both local fine details and global context is necessary for accurate predictions. Recent advancements have shown that Atrous Convolutions are crucial for improving the performance of ViTs in segmentation tasks, particularly in domains such as medical imaging. In our model, we leverage the power of ASPP and Attention mechanisms to enhance the ViT encoder’s ability to capture both local priors and global context, effectively enabling the model to handle complex, high-resolution segmentation tasks with greater accuracy.
Figure 2: LoRA and AtrousLoRA: A comparative overview with detailed insights into the Atrous Attention Module. Both LoRA and AtrousLoRA introduce a trainable encoder-decoder structure that operates in parallel with frozen pre-trained weights. (a) LoRA applies a low-rank constraint on the weight updates by factorizing them into smaller matrices (b) AtrousLoRA extends this approach by incorporating Atrous Attention Module into the bottleneck of LoRA, leveraging multi-scale dilated convolution operations for enhanced feature extraction. (c) The Atrous Attention Module features an Atrous Spatial Pyramid Pooling (ASPP) module with various dilation rates, global image-level pooling, and an attention mechanism that refines feature maps through element-wise multiplication with attention weights.
@@ III-A Overview
VesselSAM is a promptable segmentation model designed to enhance vascular structure segmentation in medical imaging. It builds upon the Segment Anything Model (SAM) framework while integrating AtrousLoRA, a novel module that combines Atrous Attention with Low-Rank Adaptation (LoRA) to improve segmentation accuracy and computational efficiency. To preserve the rich pre-trained representations of SAM, both the image encoder and prompt encoder remain frozen, while Atrous Attention and LoRA layers enhance the model’s ability to capture multi-scale contextual information and optimize training efficiency. The Atrous Attention module expands the receptive field through dilated convolutions, enabling the segmentation of both fine-grained vascular structures and broader anatomical features without increasing computational cost. Meanwhile, AtrousLoRA layers inserted within the frozen image encoder apply low-rank projections to reduce the number of trainable parameters, allowing for efficient fine-tuning while maintaining the integrity of the pre-trained backbone. The final segmentation output is generated by the mask decoder, which refines the fused embeddings from the image and prompt encoders through cross-attention mechanisms, ensuring accurate and robust segmentation performance. By leveraging AtrousLoRA, VesselSAM achieves state-of-the-art segmentation accuracy while significantly reducing computational overhead, making it well-suited for medical image segmentation tasks, particularly vascular segmentation in aortic imaging.